OpenStudy (anonymous):

how do i simplify this (x+2/3)^2-3

OpenStudy (amistre64):

multiply it out....

OpenStudy (amistre64):

is it x+(2/3)? inside there or: (x+2)/3

OpenStudy (anonymous):

no it's within the fraction

OpenStudy (amistre64):

help me out a little, cause its stll confusing.... try to space out terms like this :)

OpenStudy (anonymous):

ok ( x + 2 / 3 ) ^ 2 - 3

OpenStudy (amistre64):

...... good job...... but my plan backfired :)

OpenStudy (amistre64):

(x + 2/3 )^2 - 3 ??

OpenStudy (anonymous):

I assume it is x + (2/3), but i will give solution for both. [x + (2/3)]^2 - 3 = [x^2 + (4/3)x + (4/9)] - 3 = x^2 +(4/3)x - (23/9) [(x+2)/3]^2 - 3 = [(x+2)^2 / 9] - 3 = (x^2 + 4x + 4)/9 - (27/9) = (x^2 + 4x -23)/9